1. The potential difference between a storm cloud and the ground is 5.0 ⋅ 10! V. During a lightning flash, 3.0 coulombs of charge are transferred to the ground. (a) How much energy is transferred to the ground in this lightning flash? (b) If this energy were used to accelerate a 3500 kg truck from rest, how fast would the truck end up going?
1
Expert's answer
2020-04-06T09:25:17-0400

a)


"W=qU=(3)(5\cdot10^7)=1.5\cdot10^8\ J"

b)


"0.5mv^2=W"

"0.5(3500)v^2=1.5\cdot10^8"

"v=292\frac{m}{s}"